10. Dowsing is a form of divination used to locate hidden water wells, buried metals, gemstones, or other objects. A scientific explanation was proposed for this phenomena claiming that moving ones arm in the earths magnetic field could induce an emf along the arm of the dowser and cause muscles to twitch and the dowsing rod to move. The earths magnetic field could be slightly higher over water or other conducting materials, causing muscles to twitch over these objects (almost like a human metal detector). Experimentally investigate this phenomena. Use the casing from two old pens and place bent pieces of wire inside the casing. Do you see any affect as you walk around? Use your knowledge of electricity and magnetism to theoretically investigate this phenomena. Useful information for this question will include the action potential (voltage) required for muscle contraction which is 10 mV and the earths magnetic field (in Pittsburgh) is 55 Are you able to generate this potential in your arms by moving them in the earths magnetic field?
Answer #1

let us say the dowsing arms are with lenght L = 20 cm

for a wire moving with velocity v and making an angle \theta with the magnetic field B and by using the Faraday's law, the motional emf produced in the wire is given as

emf=vLBsin\theta

For a dowsing rod moving perpendicular to the magnetic field, sin\theta=sin90=1 so

emf=vLB=10 mV= v x 0.2m x 55 T

==> v=0.01/0.2 x 55 m/sec = 0.00091 m/sec = 1 mm/sec

so with a velocity of 1 mm/sec, it is possible to generate the required potential in the arms.

Please rate. Cheers.